# A Survey on Graph Neural Networks for Time Series: Forecasting, Classification, Imputation, and Anomaly Detection

### v1

@@ 6.2 Multivariate Time Series Classification
In essence, multivariate time series classification maintains fundamental similarities with its univariate counterpart, however, it introduces an additional layer of complexity: the necessity to capture intricate inter-time series dependencies.
@@
In both examples, the relationships between the variables, or inter-time series dependencies, can be naturally thought of as a network graph. Hence, they ideally suit the capabilities of GNNs as illustrated in forecasting section 4. As such, spatiotemporal GNNs, exemplified by those utilized in forecasting tasks [61], are conveniently adaptable for multivariate time series classification tasks. This adaptation can be achieved through the replacement of the final layer with a classification component. The unique design of these GNN architectures allows for the successful detection and capture of both inter-variable and inter-temporal dependencies. The primary aim here is to effectively distill the complexity of high-dimensional series data into a more comprehensible, yet equally expressive, representation that enables differentiation of time series into their representative classes [175, 176].
@@
The proficiency of spatiotemporal GNNs in decoding the complexities of multivariate time series is demonstrably showcased in the Raindrop architecture [46]. To classify irregularly sampled data where subsets of variables have missing values at certain timestamps, Raindrop adaptively learns a graph structure. It then dynamically interpolates missing observations within the embedding space, based on any available recorded data. This flexible approach ensures that the data representation remains both comprehensive and accurate, despite any irregularities in the sampling. Empirical studies provide evidence that Raindrop can maintain robust, high-performance classification, even in the face of such irregularities [46]. These findings further reinforce the versatility of spatiotemporal GNNs in time series classification, highlighting their effectiveness even in scenarios characterized by missing data and irregular sampling patterns.
7 GNNs for Time Series Imputation
TABLE V: Summary of graph neural networks for time series imputation. Task notation: “Out-of-sample”, “In-sample”, and “Both” refer to the types of imputation problems addressed by the approach. “Type” represents the imputation method as either deterministic or probabilistic. “Inductiveness” indicates if the method can generalize to unseen nodes. The remaining notations are shared with Table II.
Approach Year Venue Task Type Spatial Module Temporal Module Inductiveness Input Graph Learned Relations Graph Heuristics IGNNK [190] 2021 AAAI Out-of-sample Deterministic Spectral GNN - Yes R - SP, PC GACN [191] 2021 ICANN In-sample Deterministic Spatial GNN T-C No R - PC SATCN [192] 2021 arXiv Out-of-sample Deterministic Spatial GNN T-C Yes R - SP GRIN [50] 2022 ICLR Both Deterministic Spatial GNN T-R Yes R - SP SPIN [193] 2022 NIPS In-sample Deterministic Spatial GNN T-A No R - SP FUNS [194] 2022 ICDMW Out-of-sample Deterministic Spatial GNN T-R Yes R - - AGRN [195] 2022 ICONIP In-sample Deterministic Spatial GNN T-R No NR S - MATCN [196] 2022 IEEE IoT-J In-sample Deterministic Spatial GNN T-A No R - - PriSTI [51] 2023 arXiv In-sample Probabilistic Spatial GNN T-A No R - SP DGCRIN [197] 2023 KBS In-sample Deterministic Spatial GNN T-R No NR D - GARNN [198] 2023 Neurocomputing In-sample Deterministic Spatial GNN T-R No R - PC MDGCN [199] 2023 Transp. Res. Part C In-sample Deterministic Spatial GNN T-R No R - SP, PS
@@ 7 GNNs for Time Series Imputation
Time series imputation, a crucial task in numerous real-world applications, involves estimating missing or corrupted values within one or more data point sequences. Traditional time series imputation approaches have relied on statistical methodologies, such as mean imputation, spline interpolation [200], and regression models [201]. However, these methods often struggle to capture complex temporal dependencies and non-linear relationships within the data. While some deep neural network-based works, such as [202, 203, 204], have mitigated these limitations, they have not explicitly considered inter-time series dependencies. The recent emergence of graph neural networks has introduced new possibilities for time series imputation. GNN-based methods better characterize intricate spatial and temporal dependencies in time series data, making them particularly suitable for real-world scenarios arising from the increasing complexity of data. From a task perspective, GNN-based time series imputation can be broadly categorized into two types: In-sample imputation and out-of-sample imputation. The former involves filling in missing values within the given time series data, while the latter predicts missing values in disjoint sequences [50]. From a methodological perspective, GNN for time series imputation can be further divided into deterministic and probabilistic imputation. Deterministic imputation provides a single best estimate for the missing values, while probabilistic imputation accounts for the uncertainty in the imputation process and provides a distribution of possible values. In Table V, we summarize most of the related works on GNN for time series imputation to date, offering a comprehensive overview of the field and its current state of development.
7.1 In-Sample Imputation
@@
The majority of existing GNN-based methods primarily focus on in-sample time series data imputation. For instance, GACN [191] proposes to model spatial-temporal dependencies in time series data by interleaving GAT [117] and temporal convolution layers in its encoder. It then imputes the missing data by combining GAT and temporal deconvolution layers that map latent states back to original feature spaces. Similarly, SPIN [193] first embeds historical observations and sensor-level covariates to obtain initial time series representations. These are then processed by multi-layered sparse spatial-temporal attention blocks before the final imputations are obtained with a nonlinear transformation. GRIN [50] introduces the graph recurrent imputation network, where each unidirectional module consists of one spatial-temporal encoder and two different imputation executors. The spatial-temporal encoder adopted in this work combines MPNN [110] and GRU [114]. After generating the latent time series representations, the first-stage imputation fills missing values with one-step-ahead predicted values, which are then refined by a final one-layer MPNN before passing to the second-stage imputation for further processing. Similar works using bidirectional recurrent architectures include AGRN [195], DGCRIN [197], GARNN [198], and MDGCN [199], where the main differences lie in intermediate processes. For example, AGRN and DGCRIN propose different graph recurrent cells that integrate graph convolution and GRU to capture spatial-temporal relations, while GARNN involves the use of GAT and different LSTM [205] cells to compose a graph attention recurrent cell in its model architecture. MDGCN models time series as dynamic graphs and captures spatial-temporal dependencies by stacking bidirectional LSTM and graph convolution. Recently, a few research studies have explored probabilistic in-sample time series imputation, such as PriSTI [51], where the imputation has been regarded as a generation task. In PriSTI, a similar architecture of denoising diffusion probabilistic models [206] has been adopted to effectively sample the missing data with a spatial-temporal denoising network composed of attentive MPNN and temporal attention.
7.2 Out-of-Sample Imputation
@@
To date, only a few GNN-based methods fall into the category of out-of-sample imputation. Among these works, IGNNK [190] proposes an inductive GNN kriging model to recover signals for unobserved time series, such as a new variable or “virtual sensor” in a multivariate time series. In IGNNK, the training process involves masked subgraph sampling and signal reconstruction with the diffusion graph convolution network presented in [129]. Another similar work is SATCN [192], which also focuses on performing real-time time series kriging. The primary difference between these two works lies in the underlying GNN architectures, where SATCN proposes a spatial aggregation network combined with temporal convolutions to model the underlying spatial-temporal dependencies. It is worth noting that GRIN [50] can handle both in-sample and out-of-sample imputations, as well as a similar follow-up work [194].
8 Practical Applications
@@
Graph neural networks have been applied to a broad range of disciplines related to time series analysis. We categorize the mainstream applications of GNN4TS into six areas: Smart transportation, on-demand services, environment & sustainable energy, internet-of-things, and healthcare.
@@
Fraud Detection. As elucidated in the four-element fraud diamond [268], the perpetration of fraud necessitates not just the presence of incentive and rationalization, but also a significant degree of capability - often attainable only via coordinated group efforts undertaken at opportune moments. This suggests that fraud is typically committed by entities possessing sufficient capability, which can be primarily achieved through collective endeavors during suitable periods. Consequently, it is rare for fraudsters to operate in isolation [269, 270]. They also frequently demonstrate unusual temporal patterns in their activities, further supporting the necessity for sophisticated fraud detection measures [271, 272]. To this end, GNNs have been proposed to capture these complex relational and temporal dynamics inherent to fraud network activities. They have found successful applications in various domains, such as detecting frauds and anomalies in social networks [271, 272, 273], financial networks and systems [274, 275, 276, 277], and in several other sectors [270, 278, 279, 280, 281].
@@
Other Applications. Beyond the aforementioned sectors, the application of GNNs for time series analysis has also been extended to various other fields, such as finance [282], urban planning [283, 284], epidemic control [285, 286, 287], and particle physics [288]. As research in this area continues to evolve, it is anticipated that the application of GNNs will continue to expand, opening up new possibilities for data-driven decision making and system optimization.
9 Future Directions

### v3

@@ C.2 Multivariate Time Series Classification
In essence, multivariate time series classification maintains fundamental similarities with its univariate counterpart, however, it introduces an additional layer of complexity: the necessity to capture intricate inter-variable dependencies.
@@
In both examples, the relationships between the variables, or inter-variable dependencies, can be naturally thought of as a network graph. Hence, they ideally suit the capabilities of GNNs as illustrated in forecasting Sec. 4. As such, spatial-temporal GNNs, exemplified by those utilized in forecasting tasks [51], are conveniently adaptable for multivariate time series classification tasks. This adaptation can be achieved through the replacement of the final layer with a classification component. The unique design of these STGNN architectures enables the capture of both inter-temporal and inter-variable dependencies. The primary aim here is to effectively distill the complexity of high-dimensional series data into a more comprehensible, yet equally expressive, representation that enables differentiation of time series into their representative classes [151, 64].
@@
The proficiency of spatial-temporal GNNs in decoding the complexities of multivariate time series is demonstrably showcased in the Raindrop architecture [33]. To classify irregularly sampled data where subsets of variables have missing values at certain timestamps, Raindrop adaptively learns a graph structure. It then dynamically interpolates missing observations within the embedding space, based on any available recorded data. This flexible approach ensures that the data representation remains both comprehensive and accurate, despite any irregularities in the sampling. Empirical studies provide evidence that Raindrop can maintain robust, high-performance classification, even in the face of such irregularities [33]. These findings further reinforce the versatility of spatial-temporal GNNs in time series classification, highlighting their effectiveness even in scenarios characterized by missing data and irregular sampling patterns.
Appendix D GNNs for Time Series Imputation
@@
Time series imputation, a crucial task in numerous real-world applications, involves estimating missing values within one or more data point sequences. Traditional time series imputation approaches have relied on statistical methodologies, such as mean imputation, spline interpolation [171], and regression models [172]. However, these methods often struggle to capture complex temporal dependencies and non-linear relationships within the data. While some deep neural network-based works, such as [173, 244, 174], have mitigated these limitations, they have not explicitly considered inter-variable dependencies. The recent emergence of graph neural networks has introduced new possibilities for time series imputation. GNN-based methods better characterize intricate spatial and temporal dependencies in time series data, making them particularly suitable for real-world scenarios arising from the increasing complexity of data. From a task perspective, GNN-based time series imputation can be broadly categorized into two types: in-sample imputation and out-of-sample imputation. The former involves filling in missing values within the given time series data, while the latter predicts missing values in disjoint sequences [37]. From a methodological perspective, GNN for time series imputation can be further divided into deterministic and probabilistic imputation. Deterministic imputation provides a single best estimate for the missing values, while probabilistic imputation accounts for the uncertainty in the imputation process and provides a distribution of possible values. In Tab. 5, we summarize most of the related works on GNN for time series imputation to date, offering a comprehensive overview of the field and its current state of development.
D.1 In-sample Imputation
@@
The majority of existing GNN-based methods primarily focus on in-sample time series data imputation. For instance, GACN [65] proposes to model spatial-temporal dependencies in time series data by interleaving GAT [120] and temporal convolution layers in its encoder. It then imputes the missing data by combining GAT and temporal deconvolution layers that map latent states back to original feature spaces. Similarly, SPIN [67] first embeds historical observations and sensor-level covariates to obtain initial time series representations. These are then processed by multi-layered sparse spatial-temporal attention blocks before the final imputations are obtained with a nonlinear transformation. GRIN [37] introduces the graph recurrent imputation network, where each unidirectional module consists of one spatial-temporal encoder and two different imputation executors. The spatial-temporal encoder adopted in this work combines MPNN [115] and GRU [118]. After generating the latent time series representations, the first-stage imputation fills missing values with one-step-ahead predicted values, which are then refined by a final one-layer MPNN before passing to the second-stage imputation for further processing. Similar works using bidirectional recurrent architectures include AGRN [166], DGCRIN [167], GARNN [168], and MDGCN [81], where the main differences lie in intermediate processes. For example, AGRN and DGCRIN propose different graph recurrent cells that integrate graph convolution and GRU to capture spatial-temporal relations, while GARNN involves the use of GAT and different LSTM [245] cells to compose a graph attention recurrent cell in its model architecture. MDGCN models time series as dynamic graphs and captures spatial-temporal dependencies by stacking bidirectional LSTM and graph convolution. Recently, a few research studies have explored probabilistic in-sample time series imputation, such as PriSTI [38] and [170], where the imputation has been regarded as a generation task. In PriSTI, a similar architecture of denoising diffusion probabilistic models [246] has been adopted to effectively sample the missing data with a spatial-temporal denoising network composed of attentive MPNN and temporal attention. [170] adopts similar concept but with GAT [120] as the backbone denoising network.
D.2 Out-of-sample Imputation
@@
To date, only a few GNN-based methods fall into the category of out-of-sample imputation. Among these works, IGNNK [73] proposes an inductive GNN kriging model to recover signals for unobserved time series, such as a new variable or “virtual sensor” in a multivariate time series. In IGNNK, the training process involves masked subgraph sampling and signal reconstruction with the diffusion graph convolution network presented in [175]. Another similar work is SATCN [66], which also focuses on performing real-time time series kriging. The primary difference between these two works lies in the underlying GNN architectures, where SATCN proposes a spatial aggregation network combined with temporal convolutions to model the underlying spatial-temporal dependencies. INCREASE [169], on the other hand, further considers heterogeneous spatial and diverse temporal relations among the locations, lifting the performance of inductive spatio-temporal kriging. It is worth noting that GRIN [37] can handle both in-sample and out-of-sample imputations, as well as a similar follow-up work [165].
Appendix E Practical Applications
@@
Graph neural networks have been applied to a broad range of disciplines related to time series analysis. We categorize the mainstream applications of GNN4TS into seven areas: smart transportation, on-demand services, environment & sustainable energy, internet-of-things, physical systems, healthcare, and fraud detection.
@@
Physical systems. Systems of interacting objects are found in numerous scientific fields, including the simulation of n-body systems [187], particle physics [188], modeling of human motion dynamics [189], and prediction of molecular dynamics [190]. In graph-based deep learning, the objects are represented as nodes of a graph and GNNs have proven effective in modeling their complex interactions. Despite the challenges posed by physical constraints, promising performance has been achieved thanks to the incorporation of inductive biases from known physical laws [191, 192] and architectures that maintain the symmetries of the underlying system, such as equivariance to rotations and translations [193, 194, 190].
@@
TABLE VI: A summary of selected benchmark datasets grouped by task categories.
Task Category Dataset # Samples # Nodes Sampling Rate Missing Ratio Data Type References Forecasting & Imputation METR-LA [68] 34,272 207 5 minutes 8.109% Traffic Velocity [68, 48, 69, 51, 50], [52, 23, 61, 77, 109] PeMS-BAY [68] 52,116 325 5 minutes 0.003% Traffic Velocity [68, 48, 69, 51, 70], [50, 52, 60, 23, 61] PeMSD3 [302] 26,208 358 5 minutes 0.672% Traffic Volume [92, 50, 74, 44, 88], [75, 108] PeMSD4 [302] 16,992 307 5 minutes 3.182% Traffic Volume [92, 50, 49, 105, 71], [74, 44, 88, 75, 61] PeMSD7 [302] 28,224 883 5 minutes 0.452% Traffic Volume [54, 92, 50, 105, 74], [44, 88, 60, 75, 108] PeMSD8 [302] 17,856 170 5 minutes 0.696% Traffic Volume [92, 50, 49, 105, 71], [74, 44, 88, 75, 72] Xiamen [70] 44,064 95 5 minutes - Traffic Volume [70, 303, 169, 304] Beijing [305] 278,085 36 1 hour - Air Quality Index [305, 169, 306, 307], [308] NYC-Bike [309] 3.8 million 112 30 minutes - Trip Records [110, 310, 309, 311], [312, 313] NYC-Taxi [309] 28.1 million 192 30 minutes - Trip Records [110, 310, 309, 311], [312, 313] NYC-Crime [314] 31,799 - 85,899 256 1 day - Crime Incidents [314, 315] AQI [305] 8,759 437 30 minutes 25.7% Air Quality Index [305, 37, 109, 316], [67, 166] AQI-36 [317] 8,759 36 1 hour 13.2% Air Quality Index [317, 37, 67, 166] Anomaly Detection SMD [141] 608,342 38 1 minute 5.84% Server Machine [35, 84, 90, 79] [130, 80, 318, 62] SMAP [140] 562,800 25 1 minute 13.13% Spacecraft Telemetry [35, 89, 79, 130, 131], [80, 318] MSL [140] 132,046 55 1 minute 10.72% Spacecraft Telemetry [35, 89, 79, 130, 131] [80, 80, 318] SWaT [319] 925,010 51 1 second 11.97% Industrial Systems [36, 89, 57, 90, 56] [131, 80, 318, 62, 150] WADI [320] 1.4 million 127 1 second 5.99% Industrial Systems [36, 89, 90, 56], [131, 318, 62, 150] Classification UCR Archive[321] 40-24,000 1 Varies - Wide Applications [151, 63, 97] UEA Repository [322] 27-50,000 Varies 2-1345 - Wide Applications [64, 153]
@@
Fraud Detection. As elucidated in the four-element fraud diamond [323], the perpetration of fraud necessitates not just the presence of incentive and rationalization, but also a significant degree of capability - often attainable only via coordinated group efforts undertaken at opportune moments. This suggests that fraud is typically committed by entities possessing sufficient capability, which can be primarily achieved through collective endeavors during suitable periods. Consequently, it is rare for fraudsters to operate in isolation [324, 325]. They also frequently demonstrate unusual temporal patterns in their activities, further supporting the necessity for sophisticated fraud detection measures [197, 326]. To this end, GNNs have been proposed to capture these complex relational and temporal dynamics inherent to fraud network activities. They have found successful applications in various domains, such as detecting frauds and anomalies in social networks [197, 326, 327], financial networks and systems [328, 329, 330, 331], and in several other sectors [325, 332, 333, 334, 335, 336].
@@
Other Applications. Beyond the aforementioned sectors, the application of GNNs for time series analysis has also been extended to various other fields, such as finance [195], urban planning [337, 338], epidemic control [339, 340, 341, 342], recommender systems [199, 343, 344], and manufacturing [198]. As research in this area continues to evolve, it is anticipated that the application of GNN4TS will continue to expand, opening up new possibilities for data-driven decision making and system optimization.
Appendix F Datasets and Implementations
TABLE VII: A summary of open-source implementation of representative approaches.
Approach Year Task Category Github Repository DCRNN [68] 2018 Forecasting https://github.com/liyaguang/DCRNN STGCN [54] 2018 Forecasting https://github.com/VeritasYin/STGCN_IJCAI-18 ST-MetaNet [76] 2019 Forecasting https://github.com/panzheyi/ST-MetaNet ASTGCN [86] 2019 Forecasting https://github.com/guoshnBJTU/ASTGCN-2019-pytorch Graph WaveNet [48] 2019 Forecasting https://github.com/nnzhan/Graph-WaveNet MRA-BGCN [69] 2020 Forecasting https://github.com/wumingyao/MAR-BGCN_GPU_pytorch MTGNN [51] 2020 Forecasting https://github.com/nnzhan/MTGNN GMAN [70] 2020 Forecasting https://github.com/zhengchuanpan/GMAN StemGNN [50] 2020 Forecasting https://github.com/microsoft/StemGNN GTS [52] 2021 Forecasting https://github.com/chaoshangcs/GTS Z-GCNETs [71] 2021 Forecasting https://github.com/Z-GCNETs/Z-GCNETs STGODE [74] 2021 Forecasting https://github.com/square-coder/STGODE STFGNN [44] 2021 Forecasting https://github.com/MengzhangLI/STFGNN TPGNN [60] 2022 Forecasting https://github.com/zyplanet/TPGNN MTGODE [23] 2022 Forecasting https://github.com/TrustAGI-Lab/MTGODE STG-NCDE [75] 2022 Forecasting https://github.com/jeongwhanchoi/STG-NCDE STEP [61] 2022 Forecasting https://github.com/zezhishao/STEP CaST [91] 2023 Forecasting https://github.com/yutong-xia/CaST GPT-ST [110] 2023 Forecasting https://github.com/HKUDS/GPT-ST MTAD-GAT [35] 2020 Anomaly Detection https://github.com/mangushev/mtad-gat GDN[36] 2021 Anomaly Detection https://github.com/d-ailin/GDN GANF[57] 2022 Anomaly Detection https://github.com/EnyanDai/GANF VGCRN[79] 2022 Anomaly Detection https://github.com/BoChenGroup/DVGCRN STGAN[80] 2022 Anomaly Detection https://github.com/dleyan/STGAN CST-GL[62] 2023 Anomaly Detection https://github.com/huankoh/CST-GL Time2Graph+[96] 2021 Classification https://github.com/petecheng/Time2GraphPlus RainDrop[33] 2022 Classification https://github.com/mims-harvard/Raindrop SimTSC[63] 2022 Classification https://github.com/daochenzha/SimTSC TodyNet[64] 2023 Classification https://github.com/liuxz1011/TodyNet MTS2Graph[153] 2024 Classification https://github.com/raneeny/MTS2Graph IGNNK [73] 2021 Imputation https://github.com/Kaimaoge/IGNNK GRIN [37] 2022 Imputation https://github.com/Graph-Machine-Learning-Group/grin SPIN [67] 2022 Imputation https://github.com/Graph-Machine-Learning-Group/spin PriSTI [38] 2023 Imputation https://github.com/LMZZML/PriSTI INCREASE [169] 2023 Imputation https://github.com/zhengchuanpan/INCREASE
We summarize the common benchmark datasets and the open-sourced implementation of representative models, which are listed in Table VI and Table VII.
Forecasting and Imputation Benchmarks. The METR-LA [68] dataset is a prominent benchmark for traffic modeling and forecasting. It comprises data collected from 207 sensors on highways in Los Angeles, recording speed in miles per hour from March 1st to June 30th, 2012. Similarly, the PEMS-BAY [68] dataset is another widely utilized benchmark for traffic modeling and forecasting. Collected by the California State Transportation Agency (CalSTA), it includes data from 325 sensors in the Bay Area, recording speeds in miles per hour from January 1st to May 31st, 2017.
PeMS [302] datasets such as PeMSD3, PeMSD4, PeMSD7, and PeMSD8 are derived from the Caltrans Performance Measurement System (PeMS). These datasets are collected in real-time every 30 seconds and aggregated into 5-minute intervals from the raw data. They typically contain various measurements, such as vehicle count and traffic velocity; however, the mentioned four PeMS datasets focus specifically on traffic volume.
The Xiamen [70] dataset includes five months of data from 95 traffic sensors in Xiamen, China, spanning from August 1st, 2015, to December 31st, 2015, with a sampling rate of five minutes. The Beijing [305] dataset contains pollutant concentration data (PM2.5) from 36 sensors in Beijing, recorded over one year from May 1st, 2014, to April 30th, 2015.
The NYC-Taxi [309] dataset features taxi trip records from New York City (NYC) for the period from January 1st to March 1st, 2015. The NYC-Bike [309] dataset includes bike trajectory data collected from the NYC Citi Bike system between July 1st and August 29th, 2016. The NYC-Crime [314] dataset records various types of crime occurrences in NYC, such as robbery and larceny, using a 1 km $\times$ 1 km spatial grid to generate an adjacency matrix.
The AQI [305] dataset consists of recordings of several air quality indices from 437 monitoring stations across 43 Chinese cities, with a focus on the PM2.5 pollutant. The AQI-36 is a reduced version of this dataset, containing data from only 36 sensors.
Anomaly Detection Benchmarks. For time-series anomaly detection, significant attention has been directed towards a select few real-world datasets, including SMD[141], MSL[140], SMAP[140], SWAT[319] and WADI[320]. These datasets have been pivotal in driving research and development within the field. However, there have been critical investigations into their limitations [345, 145, 346], particularly regarding how well they represent real-world scenarios[347]. These studies highlight the need for more diverse and representative datasets to ensure that anomaly detection models can perform robustly in practical applications.
To address these limitations, many models are also evaluated using internal datasets from proprietary investigations, providing a broader perspective on their performance. Furthermore, there have been concerted efforts to propose and develop better datasets that more comprehensively cover various types of anomaly scenarios[237, 348]. These enhanced datasets aim to capture a wider range of anomalies and identify root-cause scenarios more effectively, thereby improving the generalizability and applicability of anomaly detection models[349].
The push towards creating more representative datasets and refining evaluation metrics underscores the ongoing efforts to enhance the reliability and real-world relevance of time-series anomaly detection. However, this also requires researchers to thoroughly benchmark their models against baseline models on these datasets in a fair and comparative manner. Comprehensive research studies that benchmark existing models are crucial for building a solid foundation for future advancements. By addressing the gaps in current benchmark datasets and ensuring models are tested against realistic scenarios, the field can develop more robust and effective solutions for detecting anomalies in diverse environments.
Classification Benchmarks. Time-series classification benchmarks, particularly in the realm of univariate and multivariate datasets, have evolved and progressed substantially. The univariate time series classification benchmarks, exemplified by the UCR archive[321], and the multivariate time series benchmarks, represented by the UEA repository[322], demonstrate considerable progress compared to other fields. This progress is attributed to dedicated efforts in carefully collecting datasets from a wide array of domains and applications, ensuring a diverse and representative sample of real-world scenarios.
These standardized benchmarks have further been extensively studied through comprehensive investigations, often referred to as Bakeoff studies[350, 351, 352]. Such studies provide a systematic framework to evaluate a wide range of methods, enabling robust comparisons and fostering methodological advancements. By rigorously assessing various approaches against a consistent set of benchmarks, researchers can identify the strengths and limitations of different models, facilitating continuous improvements in the field.
Moreover, there have been concerted efforts to refine the metrics used for model comparison, enhancing the accuracy and relevance of performance evaluations[353]. These improved metrics make the benchmark datasets more indicative of real-world performances, ensuring that the models developed and tested under these conditions are better aligned with practical applications. This focus on realistic and meaningful evaluation criteria underscores the critical role of these benchmarks in advancing the state-of-the-art in time-series classification.
Appendix G Future Directions